OpenStudy (calculusxy):

Medal will be awarded! In the expression below, each letter represents a one-digit number. Where the same letter appears, it represents the same number in each case. Each distinct letter represents a different number. In order to make the equation true, what number must replace C? AAA +AAB ABC -------- 2012

OpenStudy (calculusxy):

@amistre64 @ranga @radar @robtobey @AkashdeepDeb @phi @.Sam.

OpenStudy (calculusxy):

@skullpatrol

OpenStudy (ranga):

AAA = 100A + 10A + A = 111A AAB = 100A + 10A + B = 110A + B ABC = 100A + 10B + C Add them all up: 111A + (110A + B) + (100A + 10B + C) = 321A + 11B + C = 2012 ---- (1) What is the maximum allowed for A? 2012 / 321 = 6.27. So the maximum A can be is 6. If A were 7, 321*7 = 2247 which exceeds 2012 on the RHS of (1). So max for A is 6. Can A be anything less than 6? What if A were 5? Put A = 5 in (1): 321*5 + 11B + C = 2012 or 11B + C = 407. The maximum 11B + C can be is 99 + 9 = 108. So 11B + C cannot be 407. Therefore A cannot be anything less than 6. Put A = 6 in (1) 321*6 + 11B + C = 2012 11B + C = 86 --- (2) 86/11 = 7.8. So the maximum B can be is 7. If B falls short even by 1, the total will come down by 11 which a single digit C alone cannot compensate to add to 86 in (2). So B must be 7. A = 6, B = 7 11*7 + C = 86 C = 86 - 77 = 9 A = 6, B = 7, C = 9